# CEHR-GPT: A Scalable Multi-Task Foundation Model for Electronic Health Records

### v1

@@ Abstract
Electronic Health Records (EHRs) provide a rich, longitudinal view of patient health and hold significant potential for advancing clinical decision support, risk prediction, and data-driven healthcare research. However, most artificial intelligence (AI) models for EHRs are designed for narrow, single-purpose tasks, limiting their generalizability and utility in real-world settings. Here, we present CEHR-GPT, a general-purpose foundation model for EHR data that unifies three essential capabilities - feature representation, zero-shot prediction, and synthetic data generation - within a single architecture. To support temporal reasoning over clinical sequences, CEHR-GPT incorporates a novel time-token-based learning framework that explicitly encodes patients’ dynamic timelines into the model structure. CEHR-GPT demonstrates strong performance across all three tasks and generalizes effectively to external datasets through vocabulary expansion and fine-tuning. Its versatility enables rapid model development, cohort discovery, and patient outcome forecasting without the need for task-specific retraining.
@@ 1 Introduction
We present CEHR-GPT, a unified EHR foundation model that supports three core capabilities within a single framework: feature representation, zero-shot prediction, and synthetic data generation. By integrating all three, our model demonstrates a broader and more flexible utility compared to prior approaches.
@@
The versatility of CEHR-GPT stems from a novel time-token–based learning framework that explicitly encodes patients’ dynamic timelines into the model architecture, preserving the full temporal structure and sequence dependencies inherent in longitudinal EHR data. To our knowledge, CEHR-GPT is the first multi-task foundation model to simultaneously support all three core capabilities within a unified framework.
@@ 2 Related work
Figure 1: CEHR-GPT is built on the GPT-2 architecture with a specialized patient representation designed to capture the full longitudinal trajectory of each patient. Two additional learning objectives were applied to artificial time token embeddings to improve the modeling of temporal information, namely, time decomposition and time-to-event prediction. CEHR-GPT supports three key capabilities: patient-level feature representation, zero-shot prediction, and synthetic data generation.
@@ 3.3 Model Architecture
Our preliminary experiments supported this design choice, showing that trainable positional embeddings across different positions tended to converge to similar vectors, suggesting their limited utility in our setup. Additionally, we incorporated two specialized learning objectives, Time Decomposition (TD) and Time to Event (TTE), designed to enhance the representation of time intervals within the model. The overall architecture of the CEHR-GPT foundation model is illustrated in Figure 1.
@@
To implement this, we added a TD prediction head at the positions corresponding to ATT tokens within the output layer. Each ATT embedding is split into three sub-embeddings that predict the time components of year, month, and day. This multiclass classification scenario uses cross-entropy loss to compute the loss term, allowing the model to leverage shared TD predictions across different timescales to improve the accuracy of long-term ATT predictions. The TD loss term can be expressed as the following,
@@
The overall loss function of CEHR-GPT is defined below, which comprises the original next token prediction (NTP), denoted as $L^{ntp}_{ij}$, and the two additional learning objectives defined above,
@@ 3.7 Evaluation Setup
We evaluated CEHR-GPT for three capabilities: feature representation, zero-shot prediction, and synthetic data generation. To benchmark our model against existing approaches, we compare it with MOTOR, a SOTA EHR foundation model known for its strong performance in representation learning tasks [8]. However, since MOTOR does not support zero-shot prediction or synthetic data generation, our comparison is limited to representation performance. Furthermore, we introduce an ablation baseline—GPT—which shares the same architecture as CEHR-GPT but excludes the proposed TTE and TD learning objectives. This allows us to assess the contribution of our novel timeline design.
@@ 3.7.1 Feature Representation
Beyond linear probing, we also fine-tuned CEHR-GPT and GPT for each prediction task. The training data was split into training and validation subsets using a 90:10 ratio, with the validation set used for early stopping to prevent overfitting. We performed hyperparameter tuning using 10% of the training and validation data, running 10 trials with hyperparameters sampled uniformly: learning rates ranged from $1\times 10^{-5}$ to $5\times 10^{-5}$, and weight decay values ranged from 0.01 to 0.05. The optimal hyperparameters were selected based on the lowest validation loss. The models were then retrained on the full training set, using early stopping with a patience of 1 to monitor validation loss. MOTOR was not fine-tuned, as Steinberg et al. [8] reported that its fine-tuned and linearly probed versions exhibit comparable performance. Final performance was evaluated on the held-out test set, and we report both AUROC and AUPRC metrics. The 95% confidence intervals were computed using bootstrapping.
@@ 3.7.2 Zero-Shot Prediction
Our evaluation followed the zero-shot setup proposed by ETHOS [12], in which the patient sequence up to the prediction time was provided to CEHR-GPT to generate 50 simulated patient trajectories. Generation was terminated when either of the following conditions was met: (1) the generated tokens exceeded the defined prediction window, or (2) the outcome event(s) occurred within the prediction time. If the [END] token was encountered within the prediction window—causing premature termination—we treated the simulation as censored and discarded it, replacing it with a new one. The probability of the predicted outcome was then estimated as the proportion of simulations in which the event occurred. Using these estimated probabilities and the corresponding ground-truth labels, we computed AUROC and AUPRC scores. The standard deviation of these estimates was assessed via bootstrapping. An illustrative example of this zero-shot prediction setup is provided in Supplementary Section 9.4. Due to the computational cost associated with the generative nature of zero-shot inference, we randomly sampled 5,000 test cases from each prediction task.
@@ 3.7.4 Generalization on EHRShot Benchmark
To evaluate the generalizability of CEHR-GPT, we utilized the ehrshot benchmarking dataset, which includes 6,739 comprehensive de-identified EHR records from Stanford Medicine [9]. This dataset is structured into 15 manually created tasks—both binary and multi-class classification tasks—across three categories: Operational Outcomes (e.g., 30-day readmission), Assignment of New Diagnoses (e.g., 1-year risk of pancreatic cancer), and Anticipating Lab Test Results. Given that CEHR-GPT was specifically pre-trained on visits, conditions, procedures, and medications, we excluded the Anticipating Lab Test Results category from our analysis.
We opted for the OMOP version of ehrshot for our experiments as both CEHR-GPT and GPT were trained using the CUIMC OMOP, where medical events are coded using the standard OMOP vocabularies. For each task, we extracted all relevant events prior to the prediction time from the ehrshot OMOP and transformed them into the CEHR-GPT patient representation. We evaluated both linear probing and full fine-tuning on the downstream tasks, similar to the procedure described in Section 4.1. In both settings, we extracted the vector corresponding to the final token of the patient sequence from the CEHR-GPT output and used it to train a linear classifier. The key distinction between the two approaches is that in linear probing, the pre-trained model weights were frozen, whereas in full fine-tuning, all weights were updated during training. For linear probing, we used the original tokenizer vocabulary and excluded any events not present in it. For fine-tuning, we expanded the tokenizer vocabulary to include over 5000 standard OMOP concept IDs specific to the Stanford EHR data, which were absent in the Columbia EHR. We conducted a hyperparameter search over 10 trials, sampling learning rates uniformly between $1\times 10^{-5}$ and $1\times 10^{-4}$. Sample packing was used during training, with a maximum of 16,384 tokens allowed per batch.
@@ 4 Experiments and Results
Our model, CEHR-GPT, demonstrates versatility across three key capabilities: feature representation, zero-shot prediction, and synthetic data generation. For comparison, we include MOTOR as a SOTA EHR foundation model focused on representation learning. Additionally, we evaluate a variant of CEHR-GPT, referred to as GPT, which excludes TD and TTE learning objectives.
GPT is assessed under the same experimental conditions as CEHR-GPT for the three aforementioned capabilities, except the large-scale synthetic data generation experiment, which was excluded due to the computational cost of generating 3.7 million synthetic patients. Instead, we evaluate its long-term generation ability through a targeted analysis of the treatment pathway using a smaller synthetic cohort of 100,000 patients.
Finally, we assess the generalizability of CEHR-GPT using the ehrshot benchmarking dataset [9], which provides a diverse range of clinical prediction tasks.
@@ 4.1 Feature Representation
To evaluate CEHR-GPT’s ability to generate meaningful patient representations, we assessed its performance under both linear probing and fine-tuning settings, and compared it with MOTOR and GPT. The Receiver Operating Characteristic (ROC) curves are shown in Figure 4, while the Precision-Recall (PR) curves are provided in Supplementary Figure 6. We reported ROC-AUC and PR-AUC across all tasks in Table 3.
Under the linear probing setting, MOTOR achieved the highest AUROC and AUPRC across all tasks, outperforming both CEHR-GPT-L and GPT-L. This may reflect differences in pretraining strategies, as MOTOR is explicitly optimized for representation learning (e.g., time-to-event prediction), whereas CEHR-GPT is trained using next-token prediction.
With full model fine-tuning, both CEHR-GPT-T and GPT-T showed substantial improvements across all tasks compared to their linear probing counterparts. Both models also outperformed MOTOR on several tasks, including HF Readmission, Hospitalization, AFib Ischemic Stroke, and CAD CABG. The largest gains were observed on the AFib Ischemic Stroke task, where fine-tuning improved AUROC by 7% and AUPRC by 45%. While MOTOR demonstrated minimal performance gains from fine-tuning in prior work [8], CEHR-GPT showed competitive results, particularly when fine-tuned.
Comparing GPT and CEHR-GPT, we observed that the incorporation of time token–specific learning objectives had minimal impact on supervised prediction tasks (Table 3). However, as described in the next sections, these objectives showed its clear benefits for zero-shot generalization and synthetic data generation.
Figure 4: Receiver Operating Characteristic (ROC) curves comparing the performance across six clinical prediction tasks. Models are distinguished by unique color and line style combinations. Performance metrics show Area Under the Curve (AUC) values with 95% confidence intervals calculated using bootstrap sampling. Models are ranked in descending order of AUC performance within each subplot, with annotations positioned in the bottom right corner. The shared legend displays all evaluated models with their corresponding visual representations. Abbreviations: CEHR-GPT-L (CEHR-GPT with linear probing), CEHR-GPT-T (CEHR-GPT with fine-tuning), GPT-L (The CEHR-GPT original variant with linear probing), GPT-T (The CEHR-GPT original variant with fine-tuning), LR Baseline (Logistic Regression).
@@
Table 3: Model performance on each disease prediction task. Abbreviations: CEHR-GPT-L (CEHR-GPT with linear probing), CEHR-GPT-T (CEHR-GPT with fine-tuning), GPT-L (The CEHR-GPT original variant with linear probing), GPT-T (The CEHR-GPT original variant with fine-tuning), LR Baseline (Logistic Regression).
@@ 4.2 Zero-Shot Prediction
We evaluated the zero-shot prediction capability of CEHR-GPT using three cohorts described in Table 2, representing short-term, mid-term, and long-term risk prediction tasks: HF Readmission (30-day risk), CAD CABG (1-year risk), and T2DM HF (lifetime risk).
As shown in Table 4, CEHR-GPT achieved strong zero-shot performance across all three tasks. On the short-term task (HF Readmission), it performed comparably to GPT, while demonstrating slightly improved performance on the mid-term and long-term tasks (CAD CABG and T2DM HF, respectively). These results highlight the benefits of CEHR-GPT’s temporal modeling capabilities enabled by time tokens and its autoregressive design, which are especially important in longer-horizon prediction. Importantly, the zero-shot performance of CEHR-GPT (without task-specific fine-tuning) not only rivals the linear probing performance of pretrained models, but also outperforms count-based baselines on both HF Readmission and CAD CABG, and matches the best baseline on T2DM HF. This underscores CEHR-GPT’s ability to generalize to diverse clinical prediction tasks.
@@ 4.3 Synthetic Data Generation
Table 5 presents model performance results using synthetic data. For two tasks—Hospitalization and AFib Ischemic Stroke—models trained entirely on synthetic data performed comparably to those trained on real data when evaluated on the real test set. In contrast, performance for HF Readmission and CAD CABG was slightly lower with synthetic-only training. However, augmenting the real training set with synthetic data led to modest yet consistent improvements in test performance across all four tasks, relative to models trained on real data alone, suggesting that CEHR-GPT–generated data can serve as a useful supplement when real data is limited.
@@
Table 5: Logistic regression model performance (AUROC and AUPRC) across different datasets using synthetic datasets generated from CEHR-GPT. The columns compare the model performance when 1) trained on synthetic data alone and tested on the real test set, and 2) trained on synthetic + real data, tested on the real test set. Abbreviations: HF Readmit(HF Readmission), AFib IS (AFib Ischemic Stroke), M(Metric), R (AUROC), P (AUPRC)
In treatment pathway analyses, the synthetic dataset generated by CEHR-GPT closely replicated the cohort prevalence observed in the real-world dataset across all three cohorts, as shown in Supplementary Table 10. The Hypertension cohort showed a prevalence of 0.41% in the CEHR-GPT dataset, compared to 0.45% in the real data. Similarly, for Diabetes and Depression, the synthetic prevalences were 0.19% and 0.13%, respectively, versus 0.18% and 0.14% in the real dataset. In contrast, the GPT-generated dataset exhibited lower prevalence estimates for Hypertension (0.26%) and Diabetes (0.13%), while substantially overestimating the prevalence of Depression (0.21%). These discrepancies underscore the importance of timeline-aware modeling, demonstrating that CEHR-GPT more faithfully preserves long-term patient trajectories compared to GPT.
Privacy evaluation results are shown in Supplementary Table 11. Across all four attack types, CEHR-GPT’s synthetic dataset achieved risk scores below 0.333, a threshold recommended by Yan et al. to indicate minimal privacy risk. These results demonstrate that CEHR-GPT can generate synthetic EHR data with low measurable privacy risk while preserving clinical utility for downstream modeling.

### v2

# CEHR-XGPT: A Scalable Multi-Task Foundation Model for Electronic Health Records

@@ Abstract
Electronic Health Records (EHRs) provide a rich, longitudinal view of patient health and hold significant potential for advancing clinical decision support, risk prediction, and data-driven healthcare research. However, most artificial intelligence (AI) models for EHRs are designed for narrow, single-purpose tasks, limiting their generalizability and utility in real-world settings. Here, we present CEHR-XGPT, a general-purpose foundation model for EHR data that unifies three essential capabilities - feature representation, zero-shot prediction, and synthetic data generation - within a single architecture. To support temporal reasoning over clinical sequences, CEHR-XGPT incorporates a novel time-token-based learning framework that explicitly encodes patients’ dynamic timelines into the model structure. CEHR-XGPT demonstrates strong performance across all three tasks and generalizes effectively to external datasets through vocabulary expansion and fine-tuning. Its versatility enables rapid model development, cohort discovery, and patient outcome forecasting without the need for task-specific retraining.
@@ 1 Introduction
We present CEHR-XGPT (pronounced “seer-ex-gpt”), a unified EHR foundation model that supports three core capabilities within a single framework: feature representation, zero-shot prediction, and synthetic data generation. By integrating all three, our model demonstrates a broader and more flexible utility compared to prior approaches.
@@
The versatility of CEHR-XGPT stems from a novel time-token–based learning framework that explicitly encodes patients’ dynamic timelines into the model architecture, preserving the full temporal structure and sequence dependencies inherent in longitudinal EHR data. To our knowledge, CEHR-XGPT is the first multi-task foundation model to simultaneously support all three core capabilities within a unified framework.
@@ 2 Related work
Figure 1: CEHR-XGPT is built on the GPT-2 architecture with a specialized patient representation designed to capture the full longitudinal trajectory of each patient. Two additional learning objectives were applied to artificial time token embeddings to improve the modeling of temporal information, namely, time decomposition and time-to-event prediction. CEHR-XGPT supports three key capabilities: patient-level feature representation, zero-shot prediction, and synthetic data generation.
@@ 3.3 Model Architecture
Our preliminary experiments supported this design choice, showing that trainable positional embeddings across different positions tended to converge to similar vectors, suggesting their limited utility in our setup. Additionally, we incorporated two specialized learning objectives, Time Decomposition (TD) and Time to Event (TTE), designed to enhance the representation of time intervals within the model. The overall architecture of the CEHR-XGPT foundation model is illustrated in Figure 1.
@@
To implement this, we added a TD prediction head at the positions corresponding to ATT tokens within the output layer. Each ATT embedding is split into three sub-embeddings that predict the time components of year, month, and day. This multiclass classification scenario uses cross-entropy loss to compute the loss term, allowing the model to leverage shared TD predictions across different timescales to improve the accuracy of long-term ATT predictions. The TD loss term can be expressed as follows,
@@
The overall loss function of CEHR-XGPT is defined below, which comprises the original next token prediction (NTP), denoted as $L^{ntp}_{ij}$, and the two additional learning objectives defined above,
@@ 3.7 Evaluation Setup
We evaluated CEHR-XGPT for three capabilities: feature representation, zero-shot prediction, and synthetic data generation. To benchmark our model against existing approaches, we compare it with MOTOR, a SOTA EHR foundation model known for its strong performance in representation learning tasks [8]. However, since MOTOR does not support zero-shot prediction or synthetic data generation, our comparison is limited to representation performance. Furthermore, we introduce an ablation baseline—GPT—which shares the same architecture as CEHR-XGPT but excludes the proposed TTE and TD learning objectives. This allows us to assess the contribution of our novel timeline design.
@@ 3.7.1 Feature Representation
Beyond linear probing, we also fine-tuned CEHR-XGPT and GPT for each prediction task. The training data was split into training and validation subsets using a 90:10 ratio, with the validation set used for early stopping to prevent overfitting. We performed hyperparameter tuning using 10% of the training and validation data, running 10 trials with hyperparameters sampled uniformly: learning rates ranged from $1\times 10^{-5}$ to $5\times 10^{-5}$, and weight decay values ranged from 0.01 to 0.05. The optimal hyperparameters were selected based on the lowest validation loss. The models were then retrained on the full training set, using early stopping with a patience of 1 to monitor validation loss. MOTOR was not fine-tuned, as Steinberg et al. [8] reported that its fine-tuned and linearly probed versions exhibit comparable performance. Final performance was evaluated on the held-out test set, and we report both AUROC and AUPRC metrics. The 95% confidence intervals were computed using bootstrapping.
@@ 3.7.2 Zero-Shot Prediction
Our evaluation followed the zero-shot setup proposed by ETHOS [12], in which the patient sequence up to the prediction time was provided to CEHR-XGPT to generate 50 simulated patient trajectories. Generation was terminated when either of the following conditions was met: (1) the generated tokens exceeded the defined prediction window, or (2) the outcome event(s) occurred within the prediction time. If the [END] token was encountered within the prediction window—causing premature termination—we treated the simulation as censored and discarded it, replacing it with a new one. The probability of the predicted outcome was then estimated as the proportion of simulations in which the event occurred. Using these estimated probabilities and the corresponding ground-truth labels, we computed AUROC and AUPRC scores. The standard deviation of these estimates was assessed via bootstrapping. An illustrative example of this zero-shot prediction setup is provided in Supplementary Section 9.4. Due to the computational cost associated with the generative nature of zero-shot inference, we randomly sampled 5,000 test cases from each prediction task.
@@ 3.7.4 Generalization on EHRShot Benchmark
To evaluate the generalizability of CEHR-XGPT, we utilized the ehrshot benchmarking dataset, which includes 6,739 comprehensive de-identified EHR records from Stanford Medicine [9]. This dataset is structured into 15 manually created tasks—both binary and multi-class classification tasks—across three categories: Operational Outcomes (e.g., 30-day readmission), Assignment of New Diagnoses (e.g., 1-year risk of pancreatic cancer), and Anticipating Lab Test Results. Given that CEHR-XGPT was specifically pre-trained on visits, conditions, procedures, and medications, we excluded the Anticipating Lab Test Results category from our analysis.
We opted for the OMOP version of ehrshot for our experiments as both CEHR-XGPT and GPT were trained using the CUIMC OMOP, where medical events are coded using the standard OMOP vocabularies. For each task, we extracted all relevant events prior to the prediction time from the ehrshot OMOP and transformed them into the CEHR-XGPT patient representation. We evaluated both linear probing and full fine-tuning on the downstream tasks, similar to the procedure described in Section 4.1. In both settings, we extracted the vector corresponding to the final token of the patient sequence from the CEHR-XGPT output and used it to train a linear classifier. The key distinction between the two approaches is that in linear probing, the pre-trained model weights were frozen, whereas in full fine-tuning, all weights were updated during training. For linear probing, we used the original tokenizer vocabulary and excluded any events not present in it. For fine-tuning, we expanded the tokenizer vocabulary to include over 5000 standard OMOP concept IDs specific to the Stanford EHR data, which were absent in the Columbia EHR. We conducted a hyperparameter search over 10 trials, sampling learning rates uniformly between $1\times 10^{-5}$ and $1\times 10^{-4}$. Sample packing was used during training, with a maximum of 16,384 tokens allowed per batch.
@@ 4 Experiments and Results
Our model, CEHR-XGPT, demonstrates versatility across three key capabilities: feature representation, zero-shot prediction, and synthetic data generation. For comparison, we include MOTOR as a SOTA EHR foundation model focused on representation learning. Additionally, we evaluate a variant of CEHR-XGPT, referred to as GPT, which excludes TD and TTE learning objectives.
GPT is assessed under the same experimental conditions as CEHR-XGPT for the three aforementioned capabilities, except the large-scale synthetic data generation experiment, which was excluded due to the computational cost of generating 3.7 million synthetic patients. Instead, we evaluate its long-term generation ability through a targeted analysis of the treatment pathway using a smaller synthetic cohort of 100,000 patients.
Finally, we assess the generalizability of CEHR-XGPT using the ehrshot benchmarking dataset [9], which provides a diverse range of clinical prediction tasks.
@@ 4.1 Feature Representation
To evaluate CEHR-XGPT’s ability to generate meaningful patient representations, we assessed its performance under both linear probing and fine-tuning settings, and compared it with MOTOR and GPT. The Receiver Operating Characteristic (ROC) curves are shown in Figure 4, while the Precision-Recall (PR) curves are provided in Supplementary Figure 6. We reported ROC-AUC and PR-AUC across all tasks in Table 3.
Under the linear probing setting, MOTOR achieved the highest AUROC and AUPRC across all tasks, outperforming both CEHR-XGPT-L and GPT-L. This may reflect differences in pretraining strategies, as MOTOR is explicitly optimized for representation learning (e.g., time-to-event prediction), whereas CEHR-XGPT is trained using next-token prediction.
With full model fine-tuning, both CEHR-XGPT-T and GPT-T showed substantial improvements across all tasks compared to their linear probing counterparts. Both models also outperformed MOTOR on several tasks, including HF Readmission, Hospitalization, AFib Ischemic Stroke, and CAD CABG. The largest gains were observed on the AFib Ischemic Stroke task, where fine-tuning improved AUROC by 7% and AUPRC by 45%. While MOTOR demonstrated minimal performance gains from fine-tuning in prior work [8], CEHR-XGPT showed competitive results, particularly when fine-tuned.
Comparing GPT and CEHR-XGPT, we observed that the incorporation of time token–specific learning objectives had minimal impact on supervised prediction tasks (Table 3). However, as described in the next sections, these objectives showed its clear benefits for zero-shot generalization and synthetic data generation.
Figure 4: Receiver Operating Characteristic (ROC) curves comparing the performance across six clinical prediction tasks. Models are distinguished by unique color and line style combinations. Performance metrics show Area Under the Curve (AUC) values with 95% confidence intervals calculated using bootstrap sampling. Models are ranked in descending order of AUC performance within each subplot, with annotations positioned in the bottom right corner. The shared legend displays all evaluated models with their corresponding visual representations. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning), LR Baseline (Logistic Regression).
@@
Table 3: Model performance on each disease prediction task. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning), LR Baseline (Logistic Regression).
@@ 4.2 Zero-Shot Prediction
We evaluated the zero-shot prediction capability of CEHR-XGPT using three cohorts described in Table 2, representing short-term, mid-term, and long-term risk prediction tasks: HF Readmission (30-day risk), CAD CABG (1-year risk), and T2DM HF (lifetime risk).
As shown in Table 4, CEHR-XGPT achieved strong zero-shot performance across all three tasks. On the short-term task (HF Readmission), it performed comparably to GPT, while demonstrating slightly improved performance on the mid-term and long-term tasks (CAD CABG and T2DM HF, respectively). These results highlight the benefits of CEHR-XGPT’s temporal modeling capabilities enabled by time tokens and its autoregressive design, which are especially important in longer-horizon prediction. Importantly, the zero-shot performance of CEHR-XGPT (without task-specific fine-tuning) not only rivals the linear probing performance of pretrained models, but also outperforms count-based baselines on both HF Readmission and CAD CABG, and matches the best baseline on T2DM HF. This underscores CEHR-XGPT’s ability to generalize to diverse clinical prediction tasks.
@@ 4.3 Synthetic Data Generation
Table 5 presents model performance results using synthetic data. For two tasks—Hospitalization and AFib Ischemic Stroke—models trained entirely on synthetic data performed comparably to those trained on real data when evaluated on the real test set. In contrast, performance for HF Readmission and CAD CABG was slightly lower with synthetic-only training. However, augmenting the real training set with synthetic data led to modest yet consistent improvements in test performance across all four tasks, relative to models trained on real data alone, suggesting that CEHR-XGPT–generated data can serve as a useful supplement when real data is limited.
@@
Table 5: Logistic regression model performance (AUROC and AUPRC) across different datasets using synthetic datasets generated from CEHR-XGPT. The columns compare the model performance when 1) trained on synthetic data alone and tested on the real test set, and 2) trained on synthetic + real data, tested on the real test set. Abbreviations: HF Readmit(HF Readmission), AFib IS (AFib Ischemic Stroke), M(Metric), R (AUROC), P (AUPRC)
In treatment pathway analyses, the synthetic dataset generated by CEHR-XGPT closely matched real-world prevalence for HTN and Depression, though it underestimated Diabetes (Supplementary Table 10): HTN 0.41% vs. 0.45%, Diabetes 0.11% vs. 0.18%, and Depression 0.18% vs. 0.14%. In contrast, the GPT dataset underestimated HTN (0.26%) and Diabetes (0.10%) while only aligning on Depression (0.14%). These findings underscore the value of timeline-aware modeling, with CEHR-XGPT more faithfully preserving long-term patient trajectories than GPT.
Privacy evaluation results are shown in Supplementary Table 11. Across all four attack types, CEHR-XGPT’s synthetic dataset achieved risk scores below 0.333, a threshold recommended by Yan et al. to indicate minimal privacy risk. These results demonstrate that CEHR-XGPT can generate synthetic EHR data with low measurable privacy risk while preserving clinical utility for downstream modeling.